# Gravitational wave probes of particle dark matter: a review

### v4

@@ VI.3.3 Constraints on luminous dark photon clouds
As alluded to in Section V, if dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it would produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al. (2023a). This additional coupling implies that soon after exponential growth of the boson cloud begins, electromagnetic fields act on charged particles in the vicinity of the BH, producing electron/positron pairs essentially out of vacuum, i.e. through the photon-assisted Schwinger mechanism Sauter (1931); Heisenberg and Euler (1936); Schwinger (1951). At this point, a pair-production cascade ensues Dunne et al. (2009); Monin and Voloshin (2010), sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium. This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g. at magnetic reconnection sites, Landau damping, turbulence). The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and and also periodic Siemonsen et al. (2023a). If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to that from known pulsars in the ATNF catalog Manchester et al. (2005).
@@
The absence of observed CWs from pulsars, combined with the possibility that dark-photon clouds could act as “pulsar mimickers,” motivates searches for such clouds around sources in the ATNF catalog that might be misidentified as pulsars. Ref. Mirasola et al. (2025) explores this scenario by analyzing a subset of spinning-up “pulsars” in the catalog that could instead be stellar-mass BHs surrounded by luminous dark-photon clouds. Three analysis strategies are considered: one assuming perfect phase coherence between the electromagnetic emission and any associated GW signal; one allowing for small frequency deviations around the measured pulsar frequency; and one that is most robust to theoretical uncertainties but least sensitive to GW emission. From these methods, the authors derive upper limits on the dark-photon–photon kinetic mixing parameter as a function of boson mass for the 34 sources studied, shown in Fig. 21.
@@
To produce these bounds, assumptions are made about both the BH population and their electromagnetic signatures, so the precise results are not fully robust against variations in these choices. The adopted BH distributions include a stellar-tracing spatial distribution, a Salpeter mass function, uniform spin and age distributions, and a total galactic population of $10^{8}$ BHs. Even under these assumptions, the analysis requires only $\sim 10$ detectable “events” (i.e., BHs with GW amplitudes exceeding the strain upper limit), which is conservative compared to the $\mathcal{O}(10^{3})$ events expected at certain coupling–mass values (see Fig. 3 of Mirasola et al. (2025)). Ultimately, the study excludes dark-photon masses in the range $\sim[10^{-13},10^{-12}]$ eV with kinetic mixing larger than roughly $10^{-9},10^{-7}]$.
@@
Figure 21: Taken from Mirasola et al. (2025). For each of the radio pulsar mimickers targeted in Mirasola et al. (2025), the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits. Each vertical line corresponds to one radio pulsar, whose frequency fixes the mass of the dark photon. The color corresponds to the number of events that should have been detected, assuming a “optimistic” population model for BHs, i.e. using a larger maximum mass and spin for the assumed BH population with respect to the “pessimistic” population also constrained. The bounds reproduced from other experiments can be found in O’Hare (2020), specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi et al. (2009); Caputo et al. (2020) (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy et al. (2024) (light gray). The upper limits arising from this search surpass those from existing experiments, and indicate that between $10$ and $10^{4}$ events could have been detected at couplings of between $\sim[10^{-10},10^{-7}]$.
@@ VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada et al. (2019), the authors designed two models of the source of this stochastic GW background: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background from scalar boson clouds Tsukada et al. (2019). No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. Here, “optimistic’ means assuming rapidly spinning and frequently forming BHs in the galaxy. This limitation highlighted the degeneracy between boson mass and poorly constrained astrophysical parameters, particularly BH spins.
The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al. (2021), in particular disfavoring vector boson cloud masses between $0.8,6]\times 10^{-13}$ eV at 95% confidence. Ref. Yuan et al. (2022) has also disfavored bosons with masses between $1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
Beyond scalar and vector cases, Ref. Guo et al. (2024b) investigated tensor bosons, computing the corresponding stochastic GW background and performing a dedicated search. Their analysis disfavored boson masses between $10^{-13.4}$ and $10^{-11.7}$ eV, again with the precise range depending on the assumed range of uniform spin distribution .
@@ VI.3.5 Constraints from the spins of detected binary black holes
We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al. (2015a, 2017); Brito et al. (2017); Baryakhtar et al. (2017); Cardoso et al. (2018); Ghosh and Sachdeva (2021). Recently, a Bayesian method has been developed by combining the information from each system Ng et al. (2021b), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging. Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al. (2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al. (2021a). The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance. The study presented in Ng et al. (2021a) tentatively rules out scalar bosons in the mass range $1.3,2.7]\times 10^{-13}$ eV.
These constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson; however, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs.
Thus, these kinds of analyses complement those previously that look for CWs around isolated, rotating BHs. Both have their advantages: CW analyses described above allow us to probe unknown sources and known, nearby sources whose signals we expect to be stronger than those arising from remnants of mergers. Likewise, the spins and masses of BHs formed in binary system are well measured, and the constraints obtained in Ng et al. (2021a) will only improve as we detect more and more BHs.
@@ VI.4 Prospects for GW probes of boson clouds
At the moment, despite the vast theoretical playground of ultralight dark matter boson clouds, only a few searches have actually been performed for GWs due to boson-boson annihilations. This is primarily due to the fact that the rich physics present in boson cloud binary black hole systems, and from transitions in isolated ones, can only be probed in space-based GW detectors, which are expected to be operational in the 2030s. Moreover, scalar boson clouds are by far easier to work with theoretically than vector or tensors, which has restricted the search parameter space primarily to what can be probed by CW searches. However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen and East (2020); Siemonsen et al. (2023b), new methods are being developed to search for annihilating vector boson clouds as well Jones et al. (2023, 2025). Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond. Such known unknowns limit potential constraints to be quite model agnostic (not so astrophysically informative) or too tied to uncertain BH population models. These conundrums are not yet resolved, and probably will not be until we start detecting significantly more BHs than we currently do.
@@
The chosen targets of GW searches should also be evaluated. While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki et al. (2015b), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al. (2019), though could be seen up to Gpc now if they are composed of vector bosons Jones et al. (2023). On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen and East (2020), and so the robustness of search results by targeting remnants would have to be evaluated.
@@
All-sky searches for such boson cloud systems could also be enhanced, especially in light of boson self-interactions that could alter the GW signal significantly as discussed in Section V.3. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al. (2023, 2025), could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider the possibility that the cloud could spin-up, which, in practice, fixed a maximum spin-up to which the search was sensitive (such that the spin-up did not induce a freqeuncy shift by more than one bin $\delta f$ in $T_{\text{obs}}$, i.e. $\dot{f}T_{\text{obs}}<\delta f\equiv 1/T_{\text{FFT}}$). Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal by exploring a wider parameter space and allowing the cloud to contract over time.
@@
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof et al. (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors in Ref. Hoof et al. (2024) encourage that posteriors on mass and spin, obtained through analyses of X-ray data from BHs, be made publicly available in order to incorporate into their future work.
@@ VII GW probes of soliton dark matter
Ultralight DM in the mass range $\sim[10^{-24},10^{-22}]$ eV has a Compton wavelength comparable to the size of galactic halos Hui et al. (2017); Sin (1994). In this regime, DM can form self-gravitating structures, or solitons, at the centers of galaxies, which remain stable due to quantum degeneracy pressure Schive et al. (2014).
A natural approach to probing this mass range of DM is through pulsar timing arrays, whose frequency sensitivity aligns with the nanohertz oscillation frequencies expected from DM. In this case, solitons can exert dynamical friction on binary systems, accelerating their inspiral and suppressing the gravitational wave strain amplitude at nanohertz frequencies Aghaie et al. (2024).
Alternatively, soliton DM can be framed as a low-frequency modulation of relatively higher-frequency GW. The GW frequency is modified during its propagation through the DM halo, undergoing a gravitational redshift or ’heterodyning’, analogous to the Sachs-Wolfe effect that imprints low-frequency modulations on high-frequency photons in the cosmic microwave background Sachs and Wolfe (1967). Two CW sources have been considered as high-frequency carriers for this modulation Blas et al. (2025): (1) inspiraling white-dwarf binaries, observable by space-based GW interferometers, and (2) non-axisymmetric rotating neutron stars that spin down over time.
@@ VII.1 Constraints from pulsar timing arrays
The evidence for a GW background from NANOGrav results can be used to constrain soliton DM Aghaie et al. (2024). Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit and suppressing the GW strain at nanohertz frequencies. Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig. 1 of Afzal et al. (2023), Ref. Aghaie et al. (2024) derives constraints on the soliton DM mass-to-BH mass ratio parameter space, as shown in Fig. 22. These constraints are obtained using Eq. 17 from Aghaie et al. (2024), which models energy loss due to GW emission in the presence of ultralight dark matter. This equation incorporates astrophysical inputs such as galaxy-stellar mass functions, galaxy merger rates, and galaxy pair fractions Chen et al. (2019). The constraints are obtained by integrating Eq. 17 over these distributions, taking into account variations in the assumptions about the ultralight dark-matter mass and the population of supermassive BHs. However, these constraints are sensitive to assumptions regarding astrophysical parameters. For example, the specific values chosen for galaxy merger rates, stellar mass functions, and the fraction of galaxies with supermassive BHs affect the final upper limits.
Figure 22: Taken from Aghaie et al. (2024). Upper limits on the DM mass and soliton-to-supermassive black hole mass ratio parameter space, derived from NANOGrav data. These constraints are based on the notion that the GW background from the inspiral of supermassive BHs, for which NANOGrav has found evidence, would be modulated by the dynamical friction exerted by soliton dark matter. Specifically, the presence of soliton DM alters the supermassive BH inspiral dynamics, resulting in a suppression of the GW strain at nanohertz frequencies. The magenta lines represent two distinct sets of scaling relations derived from numerical simulations: solid lines correspond to DM-only simulations, where the soliton DM is the sole interacting component, while dashed lines incorporate the influence of supermassive BHs, suggesting a modification of the scaling relations due to the presence of these massive objects.

### v5

@@ VI.3.3 Constraints on luminous dark photon clouds
As alluded to in Section V, if dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it would produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen:2022ivj. This additional coupling implies that soon after exponential growth of the boson cloud begins, electromagnetic fields act on charged particles in the vicinity of the BH, producing electron/positron pairs essentially out of vacuum, i.e. through the photon-assisted Schwinger mechanism Sauter:1931zz; Heisenberg:1936nmg; Schwinger:1951nm. At this point, a pair-production cascade ensues Dunne:2009gi; Monin:2010qj, sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium. This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g. at magnetic reconnection sites, Landau damping, turbulence). The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and and also periodic Siemonsen:2022ivj. If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to that from known pulsars in the ATNF catalog Manchester:2004bp.
@@
The absence of observed CWs from pulsars, combined with the possibility that dark-photon clouds could act as “pulsar mimickers,” motivates searches for such clouds around sources in the ATNF catalog that might be misidentified as pulsars. Ref. Mirasola:2025car explores this scenario by analyzing a subset of spinning-up “pulsars” in the catalog that could instead be stellar-mass BHs surrounded by luminous dark-photon clouds. Three analysis strategies are considered: one assuming perfect phase coherence between the electromagnetic emission and any associated GW signal; one allowing for small frequency deviations around the measured pulsar frequency; and one that is most robust to theoretical uncertainties but least sensitive to GW emission. From these methods, the authors derive upper limits on the dark-photon–photon kinetic mixing parameter as a function of boson mass for the 34 sources studied, shown in Fig. 21.
@@
To produce these bounds, assumptions are made about both the BH population and their electromagnetic signatures, so the precise results are not fully robust against variations in these choices. The adopted BH distributions include a stellar-tracing spatial distribution, a Salpeter mass function, uniform spin and age distributions, and a total galactic population of $10^{8}$ BHs. Even under these assumptions, the analysis requires only $\sim 10$ detectable “events” (i.e., BHs with GW amplitudes exceeding the strain upper limit), which is conservative compared to the $\mathcal{O}(10^{3})$ events expected at certain coupling–mass values (see Fig. 3 of Mirasola:2025car). Ultimately, the study excludes dark-photon masses in the range $\sim[10^{-13},10^{-12}]$ eV with kinetic mixing larger than roughly $10^{-9},10^{-7}]$.
@@
Figure 21: Taken from Mirasola:2025car. For each of the radio pulsar mimickers targeted in Mirasola:2025car, the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits. Each vertical line corresponds to one radio pulsar, whose frequency fixes the mass of the dark photon. The color corresponds to the number of events that should have been detected, assuming a “optimistic” population model for BHs, i.e. using a larger maximum mass and spin for the assumed BH population with respect to the “pessimistic” population also constrained. The bounds reproduced from other experiments can be found in AxionLimits, specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi:2009iz; Caputo:2020bdy (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy:2024ozh (light gray). The upper limits arising from this search surpass those from existing experiments, and indicate that between $10$ and $10^{4}$ events could have been detected at couplings of between $\sim[10^{-10},10^{-7}]$.
@@ VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada:2018mbp, the authors designed two models of the source of this stochastic GW background: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background from scalar boson clouds Tsukada:2018mbp. No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. Here, “optimistic’ means assuming rapidly spinning and frequently forming BHs in the galaxy. This limitation highlighted the degeneracy between boson mass and poorly constrained astrophysical parameters, particularly BH spins.
The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada:2020lgt, in particular disfavoring vector boson cloud masses between $0.8,6]\times 10^{-13}$ eV at 95% confidence. Ref. Yuan:2022bem has also disfavored bosons with masses between $1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
Beyond scalar and vector cases, Ref. Guo:2023gfc investigated tensor bosons, computing the corresponding stochastic GW background and performing a dedicated search. Their analysis disfavored boson masses between $10^{-13.4}$ and $10^{-11.7}$ eV, again with the precise range depending on the assumed range of uniform spin distribution .
@@ VI.3.5 Constraints from the spins of detected binary black holes
We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki:2014wva; Arvanitaki:2016qwi; Brito:2017zvb; Baryakhtar:2017ngi; Cardoso:2018tly; Ghosh:2021zuf. Recently, a Bayesian method has been developed by combining the information from each system Ng:2019jsx, assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging. Such a study has been applied to the second GW transient catalog (GWTC-2) LIGOScientific:2020ibl, and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng:2020ruv. The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance. The study presented in Ng:2020ruv tentatively rules out scalar bosons in the mass range $1.3,2.7]\times 10^{-13}$ eV.
These constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson; however, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters:1964zz; Wen:2005xn), and are a function of the (strongest) BH mergers that we see with GWs.
Thus, these kinds of analyses complement those previously that look for CWs around isolated, rotating BHs. Both have their advantages: CW analyses described above allow us to probe unknown sources and known, nearby sources whose signals we expect to be stronger than those arising from remnants of mergers. Likewise, the spins and masses of BHs formed in binary system are well measured, and the constraints obtained in Ng:2020ruv will only improve as we detect more and more BHs.
@@ VI.4 Prospects for GW probes of boson clouds
At the moment, despite the vast theoretical playground of ultralight dark matter boson clouds, only a few searches have actually been performed for GWs due to boson-boson annihilations. This is primarily due to the fact that the rich physics present in boson cloud binary black hole systems, and from transitions in isolated ones, can only be probed in space-based GW detectors, which are expected to be operational in the 2030s. Moreover, scalar boson clouds are by far easier to work with theoretically than vector or tensors, which has restricted the search parameter space primarily to what can be probed by CW searches. However, as the theoretical calculations of vector boson cloud waveforms improve Siemonsen:2019ebd; Siemonsen:2022yyf, new methods are being developed to search for annihilating vector boson clouds as well Jones:2023fzz; Jones:2024fpg. Additionally, there are also significant astrophysical uncertainties, namely: what are the distributions of BH spins, masses, distances and ages in the galaxy and beyond. Such known unknowns limit potential constraints to be quite model agnostic (not so astrophysically informative) or too tied to uncertain BH population models. These conundrums are not yet resolved, and probably will not be until we start detecting significantly more BHs than we currently do.
@@
The chosen targets of GW searches should also be evaluated. While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki:2014faa, these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi:2018pzk, though could be seen up to Gpc now if they are composed of vector bosons Jones:2023fzz. On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen:2019ebd, and so the robustness of search results by targeting remnants would have to be evaluated.
@@
All-sky searches for such boson cloud systems could also be enhanced, especially in light of boson self-interactions that could alter the GW signal significantly as discussed in Section V.3. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones:2023fzz; Jones:2024fpg, could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider the possibility that the cloud could spin-up, which, in practice, fixed a maximum spin-up to which the search was sensitive (such that the spin-up did not induce a freqeuncy shift by more than one bin $\delta f$ in $T_{\text{obs}}$, i.e. $\dot{f}T_{\text{obs}}<\delta f\equiv 1/T_{\text{FFT}}$). Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal by exploring a wider parameter space and allowing the cloud to contract over time.
@@
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof:2024quk attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors in Ref. Hoof:2024quk encourage that posteriors on mass and spin, obtained through analyses of X-ray data from BHs, be made publicly available in order to incorporate into their future work.
@@ VII GW probes of soliton dark matter
Ultralight DM in the mass range $\sim[10^{-24},10^{-22}]$ eV has a Compton wavelength comparable to the size of galactic halos Hui:2016ltb; Sin:1992bg. In this regime, DM can form self-gravitating structures, or solitons, at the centers of galaxies, which remain stable due to quantum degeneracy pressure Schive:2014dra.
A natural approach to probing this mass range of DM is through pulsar timing arrays, whose frequency sensitivity aligns with the nanohertz oscillation frequencies expected from DM. In this case, solitons can exert dynamical friction on binary systems, accelerating their inspiral and suppressing the gravitational wave strain amplitude at nanohertz frequencies Aghaie:2023lan.
Alternatively, soliton DM can be framed as a low-frequency modulation of relatively higher-frequency GW. The GW frequency is modified during its propagation through the DM halo, undergoing a gravitational redshift or ’heterodyning’, analogous to the Sachs-Wolfe effect that imprints low-frequency modulations on high-frequency photons in the cosmic microwave background Sachs:1967er. Two CW sources have been considered as high-frequency carriers for this modulation Blas:2024duy: (1) inspiraling white-dwarf binaries, observable by space-based GW interferometers, and (2) non-axisymmetric rotating neutron stars that spin down over time.
@@ VII.1 Constraints from pulsar timing arrays
The evidence for a GW background from NANOGrav results can be used to constrain soliton DM Aghaie:2023lan. Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit and suppressing the GW strain at nanohertz frequencies. Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig. 1 of NANOGrav:2023hvm, Ref. Aghaie:2023lan derives constraints on the soliton DM mass-to-BH mass ratio parameter space, as shown in Fig. 22. These constraints are obtained using Eq. 17 from Aghaie:2023lan, which models energy loss due to GW emission in the presence of ultralight dark matter. This equation incorporates astrophysical inputs such as galaxy-stellar mass functions, galaxy merger rates, and galaxy pair fractions Chen:2018znx. The constraints are obtained by integrating Eq. 17 over these distributions, taking into account variations in the assumptions about the ultralight dark-matter mass and the population of supermassive BHs. However, these constraints are sensitive to assumptions regarding astrophysical parameters. For example, the specific values chosen for galaxy merger rates, stellar mass functions, and the fraction of galaxies with supermassive BHs affect the final upper limits.
Figure 22: Taken from Aghaie:2023lan. Upper limits on the DM mass and soliton-to-supermassive black hole mass ratio parameter space, derived from NANOGrav data. These constraints are based on the notion that the GW background from the inspiral of supermassive BHs, for which NANOGrav has found evidence, would be modulated by the dynamical friction exerted by soliton dark matter. Specifically, the presence of soliton DM alters the supermassive BH inspiral dynamics, resulting in a suppression of the GW strain at nanohertz frequencies. The magenta lines represent two distinct sets of scaling relations derived from numerical simulations: solid lines correspond to DM-only simulations, where the soliton DM is the sole interacting component, while dashed lines incorporate the influence of supermassive BHs, suggesting a modification of the scaling relations due to the presence of these massive objects.